OpenStudy (anonymous):

Which of the following is a solution to the inequality below? x^2/3^2 - y^2/4^2

OpenStudy (anonymous):

A. (-8, -2) B. (3, 1) C. (4, 0) D. (-12, 4)

OpenStudy (lgbasallote):

\[\large \frac{x^2}{3^2} - \frac{y^2}{4^2} \le 1?\]

OpenStudy (anonymous):

yes! i think i can sub the numbers in individually.. correct?

OpenStudy (anonymous):

im thinking 3,1

OpenStudy (lgbasallote):

3,1 1ill give you 1 - 1/4 = 3/4 and that is less than 1 so correct